OpenStudy (anonymous):

sow would you solve this? 3 and 5/12 - 1and 7/12

OpenStudy (anonymous):

and = + ?

OpenStudy (radar):

3 5/12 - 1 7/12= 2 17/12 - 1 7/12=1 10/12= 1 5/6